OpenStudy (anonymous):

use the quadratic formula to solve the equation to the exact value or round to two decimal places. 3x^2-x-3=0

OpenStudy (ivancsc1996):

Look

OpenStudy (anonymous):

do u know the quadratic formula?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

sorry

OpenStudy (ivancsc1996):

\[x=\frac{ -b \pm \sqrt{b ^{2} -4(a)(c)} }{ 2(a) }\]

OpenStudy (anonymous):

so just plug in the numbers?

OpenStudy (anonymous):

where ax^2 + bx + c = 0

OpenStudy (ivancsc1996):

b=-1, a=3, c=-3

OpenStudy (ivancsc1996):

Yeah

OpenStudy (anonymous):

and plug them into the formula

OpenStudy (anonymous):

ok thanks! :)

OpenStudy (ivancsc1996):

Please do tell us the answer when you finish.

OpenStudy (anonymous):

ok but it will be like a month and a half.

OpenStudy (anonymous):

maybe more.

OpenStudy (ivancsc1996):

Jajaja take your time

OpenStudy (anonymous):

1.18 or -.85

OpenStudy (anonymous):

i have a test tomorrow on quadratic functions and i just dont get ehm in general.

OpenStudy (anonymous):

*them

OpenStudy (anonymous):

then well help :D and ill work it out now, 1 sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sry i got different answers but mine r close to urs, ill show u my working :)

OpenStudy (anonymous):

kk :)

OpenStudy (anonymous):

sry i made a mistake one sec lol

OpenStudy (anonymous):

haha ok

OpenStudy (anonymous):

u might have it right :D 1 sec

OpenStudy (anonymous):

and for x^2-x=6 i got (3,2)

OpenStudy (anonymous):

ur right!

OpenStudy (anonymous):

yay lol!how do you find the vertex and symmetry of things?

OpenStudy (anonymous):

give me a question and ill show u

OpenStudy (anonymous):

y=3x^2+12x+9

OpenStudy (anonymous):

The y value for x^2-x=6 i got (3,-2)

OpenStudy (anonymous):

u got +2 for the y value, it is -2

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

so since ur asking for the symetry that means u have to draw the curve y=3x^2+12x+9

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

actually the symetry would be the equation of the line perpendicular to the tangent to the curve at the turning point, u understand that?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ive never had a symetry question like this though but i think it makes sense

OpenStudy (anonymous):

alright... :)

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

well wish me luck on my test tomorrow i guess.

OpenStudy (anonymous):

lol u want me to show u how to do it though?

OpenStudy (anonymous):

yeah u just kept typing faces. xD

OpenStudy (anonymous):

if they want u to draw it then that should be easy, but to calculate it u need to know the method and lol

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ill calculate it then draw it but u'll have to watch how much marks is the question to know exactly wat to do, that is the trick

OpenStudy (anonymous):

kay

OpenStudy (anonymous):

by the way is that an actually question? find the symetry of y=3x^2+12x+9?

OpenStudy (anonymous):

yeah right from the review.

OpenStudy (anonymous):

which is no help.

OpenStudy (anonymous):

does it have the answer? i want to know if the answer is a drawing of the curve or the equation of the line?

OpenStudy (anonymous):

because our math teacher sucks.and apparently thesymmetry is x=-2, but idk the vertex.

OpenStudy (anonymous):

ill just ask my dad cuz im being called to dinner.thanks for the help though! bye! :)

OpenStudy (anonymous):

:/ ok i was just about to do it but ok if u r going to practice more ill be here :)